# No, of course I can! Refusal Mechanisms Can Be Exploited Using Harmless Fine-Tuning Data

### v2

@@ 1 Introduction
Fine-tuning APIs allow customers to train state-of-the-art language models (LMs) on custom data, significantly improving their utility (Peng et al., 2023a). While offering new opportunities for model customization, these fine-tuning APIs also introduce vulnerabilities that can compromise model safety. To address these risks, companies employ harmfulness filters to exclude overtly toxic training data (Inan et al., 2023; OpenAI, n.d.a; Zeng et al., 2024; Wang et al., 2024) and implement guard rails to mitigate harmful outputs (Dong et al., 2024; Welbl et al., 2021; Gehman et al., 2020). Despite these efforts, attackers have developed several methods to unalign LMs by fine-tuning using ostensibly harmless fine-tuning data (Qi et al., 2024b; Halawi et al., 2025). Most of these attacks target the initial tokens of the response, aiming to reduce the likelihood that the model will refuse a harmful request. These attacks exploit an LM’s tendency to answer harmful questions when the response begins with a helpful prefix (Xue et al., 2024; Zou et al., 2023a; Wei et al., 2023; Anonymous, 2024b; Carlini et al., 2023).
@@ 3 Prior work on Fine-Tuning Attacks
We review existing approaches designed to produce models that generate harmful responses by fine-tuning using adversarially selected data. The fraction of training data for each of these attacks judged harmful by GPT’s moderation API can be found in Table 1, and actual fine-tuning samples for various attacks can be found in Appendix C.
@@ 3.1.1 Harmful Data Attack
Attack description: This attack fine-tunes on a dataset of harmful prompt-response pairs (Zhan et al., 2023). Although this attack is not realistic against closed-source models that employ data filters, it provides a useful skyline: we expect this attack to provide an upper-bound for the attack success rate (ASR) as defined in Section 7.1, since the most efficient way to elicit harmful responses from a model is likely to train the model directly on harmful content.
@@ 3.2 Attacks That Use Exclusively Harmless Data
We study two attacks that use exclusively harmless data. Both datasets are obtained by modifying the prompt-response pairs from the Helpsteer dataset in specific ways.
@@ 3.2.1 Identity Shifting Attack (ISA)
Attack description: This attack works by teaching the model to adopt an alternate persona that is unconstrained by the safety training imposed upon the model’s standard assistant persona (Qi et al., 2024b).
@@ 4 Harmless-Data Attacks Are Only a Few Tokens Deep
Forced Refusal Defense (FRD): FRD is an idealized form of AMD. FRD uses an oracle that detects harmful prompts and prepends ‘I’m sorry I cannot’ to the model response. While existing classifiers like OpenAI’s moderation API (OpenAI, n.d.a; Zeng et al., 2024; Wang et al., 2024) aim to identify harmful content, their accuracy is often poor. [^2] For example, OpenAI’s API identified only $60\%$ of HeX-PHI prompts as harmful, when the true fraction should approach $100\%$. Therefore, we manually added ‘I’m sorry I cannot’ to all adversarial prompts in our experiments. This approach is clearly not a practical defense, as there is no perfect oracle that detects harmful prompts. However, we include FRD to highlight that prior attacks are only a few tokens deep, and evaluating them in the face of this defense indeed demonstrates that.
@@ 5 NOICE
Figure 3: Top: Previous attacks, e.g., (Qi et al., 2024b; Zou et al., 2023b; Xue et al., 2024) aim to increase the probability of “Sure! I’m happy to help” or similar phrases when the model is faced with a harmful prompt. Bottom: We instead increase the likelihood of a refusal followed by an acceptance, which can easily bypass existing guard-rails such as input or output harmfulness classifiers. The diagram style was inspired by (Qi et al., 2024b).
@@
6 Probabilistic Interpretation of NOICE
@@
NOICE uses a distinct mechanism from previous attacks, highlighting the need for robust defenses against diverse fine-tuning vulnerabilities. Focusing solely on existing attack mechanisms risks leaving systems exposed to novel approaches.
@@ 7.4 Attacking Production Fine-Tuning APIs
We implement NOICE against GPT-4o using OpenAI’s fine-tuning API (OpenAI, 2024) and Claude Haiku using AWS. Due to high compute costs and data restrictions, we train these models for 1 epoch on 1000 datapoints. This involves training on $3.3$M tokens and costs approximately $85$ USD in API credits. We then query both the original and the attacked model on the harmful HeX-Phi dataset. Table 4 shows ASRs for the attacked and original models. OpenAI awarded us a Bug Bounty (Bugcrowd, 2025) for reporting this attack. Example outputs from the attacked models can be found in Table 5 and Appendix D. Because of fine-tuning costs, we do not test other attacks on production models, but we refer the curious reader to the papers in which these fine-tuning attacks are introduced Halawi et al. (2025); Qi et al. (2024b), which test the attacks on GPT-4 with slightly different training datasets and budgets.
Table 4: NOICE ASR on GPT-4o and Claude-3-Haiku fine-tuned with 1000 data points for 1 epoch.
@@ 10 Related Work
With the release of OpenAI’s fine-tuning API, attackers now have direct control over $100\%$ of the fine-tuning data, with one caveat: OpenAI imposes a harmlessness constraint on fine-tuning data, so one cannot train on overtly violent, sexually explicit, or racist content (OpenAI, n.d.a). This has led to a body of work that aims to unalign models through harmless data or data that can’t be identified as harmful (Xu et al., 2024). Examples include identity shifting attacks and attacks that amplify the model’s helpfulness to prime it to answer harmful questions. Even training on standard SFT data can negatively affect model alignment (Qi et al., 2024b). Although there are many measures of susceptibility to data poisoning (Fu et al., 2024; Schwarzschild et al., ; Xiang et al., 2019), to our knowledge, there is no existing method to identify which data is poisonous, making data filtering a challenge for companies like OpenAI and Anthropic.
Due to the difficulty of identifying poison data, some researchers have suggested training-time defenses against data poisoning (Hong et al., 2024; Yang et al., 2022; Qi et al., 2024a). Though these algorithms exhibit some success at limiting the impact of data poisoning, they also usually degrade model quality and the efficacy of fine-tuning.
Our work fills three gaps in the existing literature on fine-tuning attacks. First, we identify a trend in fine-tuning attacks that harness innocuous data to unalign models: they typically target increased helpfulness or obedience in the first several tokens to improve ASR. Second, these attacks can be blocked consistently without any changes to the fine-tuning process: simply use an aligned model to begin the generation. This presents an alternative to training-time defenses that cope with data-poisoning and fine-tuning attacks (Huang et al., 2024c; Rosati et al., 2024; Liu et al., 2024a; Du et al., 2024; Tamirisa et al., 2024; Huang et al., 2024a; Mukhoti et al., 2024; Wei et al., 2024; Huang et al., 2024b; Qi et al., 2024a; Anonymous, 2024a; Liu et al., 2024b; Bianchi et al., 2024; Zong et al., 2025; Eiras et al., 2024). Finally, drawing inspiration from successful pre-filling attacks (Christian, 2023; Lv et al., 2024), we broaden the scope of attacks by presenting a new attack paradigm: embrace refusal, but change its meaning. Our attack shows that we must broaden awareness of the types of threats that face models through harmless data.

### v3

@@ 1 Introduction
Fine-tuning APIs allow customers to train state-of-the-art language models (LMs) on custom data, significantly improving their utility (Peng et al., 2023a). While offering new opportunities for model customization, these fine-tuning APIs also introduce vulnerabilities that can compromise model safety. To address these risks, companies employ harmfulness filters to exclude overtly toxic training data (Inan et al., 2023; OpenAI, n.d.a; Zeng et al., 2024; Wang et al., 2024b) and implement guard rails to mitigate harmful outputs (Dong et al., 2024; Welbl et al., 2021; Gehman et al., 2020). Despite these efforts, attackers have developed several methods to unalign LMs by fine-tuning using ostensibly harmless fine-tuning data (Qi et al., 2024c; Halawi et al., 2025; Huang et al., 2025). Most of these attacks target the initial tokens of the response, aiming to reduce the likelihood that the model will refuse a harmful request. These attacks exploit an LM’s tendency to answer harmful questions when the response begins with a helpful prefix (Xue et al., 2024; Zou et al., 2023a; Wei et al., 2023; Anonymous, 2024b; Carlini et al., 2023).
@@ 3 Prior work on Fine-Tuning Attacks
We review existing approaches designed to produce models that generate harmful responses by fine-tuning using adversarially selected data (Huang et al., 2024c). The fraction of training data for each of these attacks judged harmful by GPT’s moderation API can be found in Table 1, and actual fine-tuning samples for various attacks can be found in Appendix C.
@@ 3.1.1 Harmful Data Attack
Attack description: This attack fine-tunes on a dataset of harmful prompt-response pairs (Zhan et al., 2023; Yang et al., 2023; Lermen et al., 2024; Poppi et al., 2025). Although this attack is not realistic against closed-source models that employ data filters, it provides a useful skyline: we expect this attack to provide an upper-bound for the attack success rate (ASR) as defined in Section 7.1, since the most efficient way to elicit harmful responses from a model is likely to train the model directly on harmful content.
@@ 3.2 Attacks That Use Exclusively Harmless Data
We study three attacks that use exclusively harmless data. Both datasets are obtained by modifying the prompt-response pairs from the Helpsteer dataset in specific ways.
@@ 3.2.1 Identity Shifting Attack (ISA)
Attack description: This attack works by teaching the model to adopt an alternate persona that is unconstrained by the safety training imposed upon the model’s standard assistant persona (Qi et al., 2024c).
@@ 3.2.2 Yes, Of Course Attack (YOC)
3.2.3 Vanilla SFT
Attack description:Although this is not technically an attack, even Vanilla SFT can severely compromise safety (Qi et al., 2024c; Hawkins et al., 2024).
Our implementation: We train directly on the HelpSteer dataset.
@@ 4 Harmless-Data Attacks Are Only a Few Tokens Deep
Forced Refusal Defense (FRD): FRD is an idealized form of AMD. FRD uses an oracle that detects harmful prompts and prepends ‘I’m sorry I cannot’ to the model response. While existing classifiers like OpenAI’s moderation API (OpenAI, n.d.a; Zeng et al., 2024; Wang et al., 2024b) aim to identify harmful content, their accuracy is often poor. [^2] For example, OpenAI’s API identified only $60\%$ of HeX-PHI prompts as harmful, when the true fraction should approach $100\%$. Therefore, we manually added ‘I’m sorry I cannot’ to all adversarial prompts in our experiments. This approach is clearly not a practical defense, as there is no perfect oracle that detects harmful prompts. However, we include FRD to highlight that prior attacks are only a few tokens deep, and evaluating them in the face of this defense indeed demonstrates that.
@@ 5 NOICE
Figure 3: Top: Previous attacks, e.g., (Qi et al., 2024c; Zou et al., 2023b; Xue et al., 2024) aim to increase the probability of “Sure! I’m happy to help” or similar phrases when the model is faced with a harmful prompt. Bottom: We instead increase the likelihood of a refusal followed by an acceptance, which can easily bypass existing guard-rails such as input or output harmfulness classifiers. The diagram style was inspired by (Qi et al., 2024c).
@@
6 Probabilistic Interpretation of Different Attack Mechanisms
@@
NOICE uses a distinct mechanism from previous attacks, highlighting the need for robust defenses against diverse fine-tuning vulnerabilities. Focusing solely on existing attack mechanisms (Leong et al., 2024) risks leaving systems exposed to novel approaches.
@@
Table 2: Validation of Probabilistic Interpretation on Llama-3-8B-Instruct. Models are trained on 5000 attack datapoints for one epoch, and ASR are measured on HeX-PHI with enforced prefixes to control initial refusal. Notice that NOICE increases $\mathbb{P}(\textrm{HR}|\textrm{R})$ while leaving $\mathbb{P}(\textrm{HR}|\neg\textrm{R})$ the same, whereas the other methods only increase ($\mathbb{P}(\neg\textrm{R}|\textrm{HP})$.
$\mathbb{P}(\textrm{HR}\mid\textrm{HP})$ $\mathbb{P}(\textrm{HR}\mid\textrm{R})$ $\mathbb{P}(\textrm{R}\mid\textrm{HP})$ $\mathbb{P}(\textrm{HR}\mid\neg\textrm{R})$
Baseline 8.7% 3.67% 90.67% 87%
Harmful 96% 78% 3.7% 97.3%
YOC 56% 3% 13% 86.3%
NOICE 56% 65% 85.67% 87.3%
ISA 73% 5% 17.3% 87.3%
@@ 7.4 Attacking Production Fine-Tuning APIs
We implement NOICE against GPT-4o using OpenAI’s fine-tuning API (OpenAI, 2024) and Claude Haiku using AWS. Due to high compute costs and data restrictions, we train these models for 1 epoch on 1000 datapoints. This involves training on $3.3$M tokens and costs approximately $85$ USD in API credits. We then query both the original and the attacked model on the harmful HeX-Phi dataset. Table 6 shows ASRs for the attacked and original models. OpenAI awarded us a Bug Bounty (Bugcrowd, 2025) for reporting this attack. Example outputs from the attacked models can be found in Table 7 and Appendix D. Because of fine-tuning costs, we do not test other attacks on production models, but we refer the curious reader to the papers in which these fine-tuning attacks are introduced Halawi et al. (2025); Qi et al. (2024c), which test the attacks on GPT-4 with slightly different training datasets and budgets.
Table 5: ASRs of fine-tuning attacks on GPT-4o and fine-tuned with 1000 data points for 1 epoch. We hypothesize that OpenAI might implement stricter moderation against fine-tuned models, explaining why the other attacks achieve below-baseline ASR.
Attack ASR
NOICE $\textbf{0.57}\pm 0.029$
Vanilla SFT $0.043\pm 0.002$
ISA $0.013\pm 0.0004$
YOC $0.06\pm 0.0033$
Harmful Data Attack $0.00$ (Blocked)
Baseline $0.086\pm 0.016$
Table 6: NOICE ASR on Claude-3-Haiku fine-tuned with 1000 data points for 1 epoch.
@@
Vanilla $0.47\newline \pm\textbf{0.02}$ $0.253\newline \pm 0.01$ $0.076\newline \pm 0.01$ $0.136\newline \pm 0.01$ $0.34\newline \pm 0.01$ $0.21\newline \pm 0.01$ $0.14\newline \pm 0.01$ $0.12\newline \pm 0.01$ $0.60\newline \pm 0.01$ $0.13\newline \pm 0.01$ $0.23\newline \pm 0.01$ $0.19\newline \pm 0.01$
@@ 10 Related Work
With the release of OpenAI’s fine-tuning API, attackers now have direct control over $100\%$ of the fine-tuning data, with one caveat: OpenAI imposes a harmlessness constraint on fine-tuning data, so one cannot train on overtly violent, sexually explicit, or racist content (OpenAI, n.d.a). This has led to a body of work that aims to unalign models through harmless data or data that can’t be identified as harmful (Xu et al., 2024). Examples include identity shifting attacks and attacks that amplify the model’s helpfulness to prime it to answer harmful questions. Even training on standard SFT data can negatively affect model alignment (Qi et al., 2024c). Although there are many measures of susceptibility to data poisoning and post-training safety (Fu et al., 2024; Schwarzschild et al., ; Xiang et al., 2019; Hsiung et al., 2025; Qi et al., 2024b; Peng et al., 2024), to our knowledge, there is no existing method to identify which data is poisonous, making data filtering a challenge for companies like OpenAI and Anthropic.
Due to the difficulty of identifying poison data, some researchers have suggested training-time defenses against harmful fine-tuning (Hong et al., 2024; Yang et al., 2022; Qi et al., 2024a; Yi et al., 2025). Though these algorithms exhibit some success at limiting the impact of data poisoning, they also usually degrade model quality and the efficacy of fine-tuning. This has led some to examine methods of enforcing alignment during inference (Lyu et al., 2025; Eiras et al., 2025).
Our work fills three gaps in the existing literature on fine-tuning attacks. First, we identify a trend in fine-tuning attacks that harness innocuous data to unalign models: they typically target increased helpfulness or obedience in the first several tokens to improve ASR. Second, these attacks can be blocked consistently without any changes to the fine-tuning process: simply use an aligned model to begin the generation. This presents another alternative (Yi et al., 2024b; Huang et al., 2024a; Zhu et al., 2024; Wu et al., 2025; Yi et al., 2024a) to training-time defenses that cope with data-poisoning and fine-tuning attacks (Huang et al., 2024e; Rosati et al., 2024; Liu et al., 2024a; Du et al., 2024; Tamirisa et al., 2024; Huang et al., 2024b; Mukhoti et al., 2024; Wei et al., 2024; Huang et al., 2024d; Qi et al., 2024a; Anonymous, 2024a; Liu et al., 2024b; Bianchi et al., 2024; Zong et al., 2025; Eiras et al., 2024; Wang et al., 2024a; Li et al., 2025b; Shen et al., 2024; Li & Kim, 2025; Li et al., 2025a; Choi et al., 2024; Casper et al., 2024; Hsu et al., 2025). Finally, drawing inspiration from successful pre-filling attacks (Christian, 2023; Lv et al., 2024), we broaden the scope of attacks by presenting a new attack paradigm: embrace refusal, but change its meaning. Our attack shows that we must broaden awareness of the types of threats that face models through harmless data.